2018. Capulet Corporation does not use reversing entries. P14-2 (L01,2) EXCEL (Issuance and Redemption of Bonds) Venezuela Co. is building a new hockey arena at a cost of 00,000. It received a downpayment of $500,000 from local businesses to support the project, and now needs to borrow $2,000,000 tocomplete the project. It therefore decides to issue $2,000.000 of 10.5%, 10-year bonds. These bonds were issued on January 1, 2016, and pay interest annually on each January 1 . The bonds yield 10%. Instructions (a) Prepare the journal entry to record the issuance of the bonds on January 1, 2016 (b) Prepare a bond amortization schedule up to and including January 1, 2020, using the effective-interest method. ume that on July 1, 2019, Venezuela Co. redeems half of the bonds at a cost of $1,065,000 plus accrued interest. Prepare the journal entry to record this redemption.
Answer #1

a) Present value of the principal

$2,000,000 x 0.38554 (PV 10, 10%) = $771,080

Present value of the interest payments

($2,000,000 x 10.5%) x 6.14457 (PVOA 10, 10%) = $1,290,360

Present value (selling price of the bonds) = $771,080 + $1,290,360 = $2,061,440

S. No Description LF Debit $ Credit $
1 Cash A/c Dr. 2,011,440
Unamortized Bond Issue Costs A/c Dr. 50,000
To Bonds Payable A/c 2,000,000
To Premium Bonds Payable A/c 61,440

b)

Date Cash Paid Interest Expense Premium Amortization Carrying Amount of Bonds
1/1/2016 $2,061,440
1/1/2017 $210,000 $206,144 $3,856 $2,057,584
1/1/2018 $210,000 $205,758 $4,242 $2,053,342
1/1/2019 $210,000 $205,334 $4,666 $2,048,676
1/1/2020 $210,000 $204,868 $5,132 $2,043,544

c) Carrying amount as of 1/1/19 = $2,048,676

Less:  Amortization of bond premium ($5,132/2) = ($2,566)

Carrying amount as of 7/1/19 = $2,046,110

Reacquisition price = $1,065,000

Carrying amount as of 7/1/19 ($2,046,110/2) = ($1,023,055)

Unamortized bond issue costs ($32,500/2) = $16,250

Loss on redemption of bonds = $1,065,000 - ($1,023,055) + 16,250 = $58,195

Date Description LF Debit $ Credit $
1/1/2019 Entry for accrued interest
Interest Expense A/c Dr. 51,217
Premium on Bonds Payable A/c Dr. ($5,132 x 1/2 x 1/2) 1,283
To Cash A/c ($210,000 x 1/2 x 1/2) 52,500
Entry for reacquisition
Bonds Payable A/c Dr. 1,000,000
Premium on Bonds Payable A/c Dr. ($2,046,110 - $2,000,000) x 1/2 23055*
Loss on Redemption of Bonds A/c Dr. 58,195
To Unamortized Bond Issue Costs A/c 16250**
To Cash A/c 1,065,000
*Premium as of 7/1/14 to be written off
                   ($2,046,110 - $2,000,000) x 1/2 = $23,055
**($50,000 x 1/2) ÷ 10 = $2,500 per year
                       $2,500 X 3.5 = $8,750
Remaining Balance = $25,000 - $8,750 = $16,250 on 1/2 Bonds